The equation of refraction at a spherical surface is


 


μ2 / v – μ1 /u = μ2- μ1 /R


 


taking R=∞, show that this equation leads to equation


 


real depth/ apparent depth = μ21


Grade:12

1 Answers

Chetan Mandayam Nayakar
13 years ago

Dear Nikhil, 

The liquid surface is flat. This implies that it is like a sphere of infinite radius. Thus, R=∞. real depth= distance of actual object from surface(inside liquid)=u, apparent depth=distance of image from surface, within liquid=v, therefore solving the equation μ2 / v – μ1 /u = μ2- μ1 /R, taking R=infinity, we get the required result.

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar respectively : Click here to download the toolbar.. 

Askiitians Expert

Chetan Mandayam Nayakar – IIT Madras